OpenStudy (anonymous):

what is the derivative of following dx/dt=? dy/dt=? dz/dt=? x=sine^t y=cose^t z=Underoot3 e^t

OpenStudy (anonymous):

\[z=\sqrt{3}e ^{t}\]

OpenStudy (anonymous):

sin^t of which variaable ??

OpenStudy (anonymous):

given is t dear

OpenStudy (anonymous):

i wrote the question same which is written

OpenStudy (anonymous):

that is sin e^t

OpenStudy (anonymous):

oh its easy ...... the answers are as follows : dy/dt is \[e^{t}\cos(e^{t}) \] dy/dt = \[-e^{t} \sin (e^{t}) \] dz/dt = \[\sqrt{3}e^{t}\]

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

wlcme

OpenStudy (anonymous):

in the future, when you dealw ith these partial differentiations, you need to specify that it is one